OpenStudy (anonymous):

Two capacitors, one 7.0 micro-farad & the other 15 micro-farad , are connected in series across a 15 V battery. Find the charge stored on each capacitors.

OpenStudy (anonymous):

Well i do hope you know that when capacitors are connected in series, they store exactly the same amount of charge.. i can prove it.. but.. the proof must be available in your textbook :P.. so anyways.. find out the equivalent capacitance.. nd 15V comes across this equivalent capacitance, and now using the basic capacitor formula, calculate the charge stored by it..

OpenStudy (kropot72):

The joint capacitance = product * sum = (7 * 15)(7 + 15) = 4.773 micro-farads. Q = CV = 4.773 * 15 micro-coulombs. This is the charge stored on each capacitor.

OpenStudy (anonymous):

Thanks!

OpenStudy (kropot72):

You're welcome :)